OpenStudy (anonymous):

One of Madison's hobbies is to locate hidden geocaches. To find yesterday's cache, Madison was instructed to start at a water fountain at a local park. Then, the directions had her walk 30 yards due west and 16 yards due south. If this was the location of the cache, how far was it from the water fountain where she started? A. 34 yards B. 14 yards C. 23 yards D. 46 yards

OpenStudy (anonymous):

@hereweg0 plz help

OpenStudy (anonymous):

@djcool31 plz help

OpenStudy (radar):



OpenStudy (anonymous):

d

OpenStudy (radar):

How did you get 46? Sum up the square of 30 and the square of 16 take the square root of that sum. square of 30 is 30 times 30 square of 16 is 16 times 16.

OpenStudy (radar):

It is a right triangle, you have been given the length of each leg, now you only need to calculate the length of the hypotenuse.

OpenStudy (anonymous):

owe my fault b

OpenStudy (radar):

Are you using a calculator?

OpenStudy (radar):

Are you trying to learn, or just looking for answers?

OpenStudy (whpalmer4):

30*30= 16*16= Add those two numbers. Find a number in the answer choices which when multiplied with itself gives you the sum of 30*30 + 16*16

OpenStudy (radar):

Very good @whpalmer4 a calculator would not even be needed using that method.

OpenStudy (anonymous):

a

OpenStudy (anonymous):

thanks

OpenStudy (whpalmer4):

Yes, a is correct!

OpenStudy (anonymous):

bye thanks for the help